# Traveling supersolid stripe patterns in spin-orbit-coupled Bose-Einstein condensates

### v1

@@ Abstract
We consider a traveling supersolid stripe pattern in a spin-orbit-coupled Bose gas. This configuration is associated with an unequal occupation of the two single-particle energy minima, giving rise to a chemical potential difference that sets the fringe velocity. Unlike stationary stripes, the moving pattern is spin-polarized, with decreasing contrast as momentum increases, eventually leading to stripe melting and transition to the uniform plane-wave phase. The Bogoliubov spectrum of the moving stripes exhibits asymmetry under inversion of the excitation quasimomentum. At high momentum, we identify energetic and dynamical instabilities in the spin-phonon mode which transforms to the roton mode of the plane-wave phase as the stripe structure vanishes.
@@ I Introduction
In recent years, the interest in supersolidity, where superfluidity coexists with a crystal-like structure, has grown significantly (see reviews in [1, 2, 3, 4, 5, 6]). First discussed in the 1950s [7, 8, 9] and further developed in subsequent decades [10, 11, 12, 13, 14, 16, 15, 17, 18], the supersolidity has only recently been observed in ultracold boson systems inside optical resonators [19], with spin-orbit coupling [20, 21, 22], and with dipolar interactions [23, 24, 25, 26, 27], as well as in exciton-polariton systems [28, 29]. These advances sparked studies on several aspects of supersolids, including their ability to sustain a dissipationless flow, i.e., a supercurrent. At low flow velocities, the superfluid fraction–bounded strictly below one due to crystalline order [13, 18]–is inferred from the system’s response to a translation [30, 31, 32, 33, 34, 35, 36] or rotation [39, 37, 40, 38, 41] constraint. However, accessing this quantity in experiments has proven challenging [42, 43], and only very recently a measurement has been performed [44]. At higher flow velocities, supercurrent-carrying supersolids can become unstable [45, 46, 32, 47], and in rotating configurations this can lead to the formation of vortices [17, 48, 49, 50], as confirmed in recent experiments [51, 52]. An anomalous Doppler effect has also been predicted [53]. Interestingly, although these results indicate that supersolids are (under certain conditions) able to sustain a supercurrent, frictionless motion of external objects remains impossible [17, 54].
@@
A common method to generate a supercurrent in a Bose-Einstein condensate (BEC) is to impose a phase twist on its time-independent order parameter [55]. In supersolids, this leads to a configuration where the lattice associated with the normal component remains at rest, while the superfluid background flows at a velocity determined by the twist angle. An alternative approach involves considering a moving lattice pattern with velocity $\bm{v}$, making the time-dependent order parameter a function of $\bm{r}-\bm{v}t$. In Galilean-invariant systems, such as dipolar gases, these two descriptions are physically equivalent, being connected by a Galilean transformation. However, spin-orbit-coupled BECs lack Galilean invariance, a feature that profoundly alters their superfluid behavior even in the non-supersolid phases. Early studies have shown that in these phases the superfluid fraction is strictly less than unity even at zero temperature [56, 57, 58]. In addition, the critical velocity for supercurrent stability differs from the threshold for frictionless impurity motion [59, 60], and current-carrying configurations can exhibit both energetic and dynamical instabilities [61], unlike in standard condensates, where only energetic (Landau) instabilities are present [62, 63]. In the supersolid stripe phase, the combined lack of translational and Galilean invariance further reduces the superfluid density compared to the Leggett bound [58] and leads to peculiar stability conditions for current-carrying states [64].
@@
The Raman lasers responsible for generating spin-orbit coupling play a crucial role in the construction of supercurrent states within the stripe phase. In the analysis of Ref. [64], the current is generated by imposing twisted boundary conditions on the condensate order parameter. This procedure effectively constrains the system’s kinetic momentum and describes a scenario in which the superfluid background flows relative to the Raman laser frame, while the density modulations remain stationary. This is consistent with the idea that the Raman lasers pin the normal component [56]. In the rest frame of the superfluid, both the fringes and the lasers appear to move at the same velocity. These configurations correspond to the two contrasting, but physically equivalent, pictures discussed earlier: background flow versus fringe motion.
In this work, we unveil a third scenario, the one in which the density modulations and the Raman lasers move relative to each other. This relative motion renders the configuration physically distinct from the two previously discussed cases. Here, the relevant conserved quantity is the canonical momentum, which, unlike the kinetic momentum, commutes with the spin-orbit Hamiltonian [61]. In contrast to earlier scenarios, these traveling stripe patterns do not allow a simple interpretation in terms of independent motion of the superfluid or normal components. Instead, they can be understood as a generalization of the notion of an imbalanced BEC mixture to the context of spin-orbit coupling, where the relevant degrees of freedom are dressed spin states rather than bare spin components. They represent intermediate structures that smoothly interpolate between the stationary stripe phase, i.e., a balanced mixture of two dressed spin states, and the uniform plane-wave phase corresponding to a fully polarized configuration in dressed spin. Within this framework, the fringe motion arises from a population imbalance between the two dressed spin components entering the condensate order parameter. The resulting chemical potential difference drives the density fringes to propagate at a constant velocity.
The motion of density fringes relative to stationary Raman lasers is well understood in the low-velocity limit, where it corresponds to the crystal Goldstone mode of the stripe phase and can be excited by releasing a weak spin perturbation [65, 66]. The goal of this work is to systematically investigate the properties of traveling stripe configurations beyond this regime. We show that at finite momentum the moving stripes develop asymmetric density profiles in the two spin components. Moreover, they display finite spin polarization as well as reduced contrast and higher energy compared to stationary stripes, effects that intensify with increasing momentum. Eventually, the fringes disappear, and the system undergoes a transition to a uniform plane-wave condensate. An analysis of the excitation spectrum of the moving supersolid reveals both energetic and dynamical instabilities emerging at sufficiently high momentum, involving the spin phonon branch. This mode becomes the mode with the roton minimum of the plane-wave phase [67].
@@
This paper is structured as follows. After briefly reviewing the model for a BEC with Raman-induced spin-orbit coupling (Sec. II), we outline the construction of traveling stripe solutions (Sec. III). These solutions are then derived and analyzed in detail (Sec. IV), with numerical results compared with perturbative analytical estimates. Section V focuses on the Bogoliubov spectrum and the instabilities of the moving stripes. Conclusions are presented in Sec. VI. Technical details of the perturbative approach and full expressions of relevant coefficients appearing in the formulas of the main text are given in Appendices A and B, respectively.
@@ II The model
where $\bm{p}=-\mathrm{i}\hbar\nabla_{\bm{r}}$ is the canonical momentum operator. The spin-orbit coupling strength is proportional to the momentum imparted by the Raman lasers, which is $-2\hbar k_{R}\hat{\bm{e}}_{x}$, with $\hat{\bm{e}}_{x}$ the unit vector along the $x$ axis. The corresponding energy scale is $E_{R}=\hbar^{2}k_{R}^{2}/2m$. The parameter $\Omega_{R}$ quantifies the Raman coupling strength, while $\delta_{R}$ represents the Raman detuning, which we shall set to zero from now on.
@@
is conserved. We note that $\bm{p}$ differs from the kinetic momentum $\bm{P}=\bm{p}-\hbar k_{R}\sigma_{z}\hat{\bm{e}}_{x}$, that is, the operator whose expectation value, $\langle\bm{P}\rangle=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\bm{P}\Psi$, gives the center-of-mass velocity of the system multiplied by the total mass $Nm$.
@@
Due to the term $-\hbar k_{R}\sigma_{z}$ arising from the spin-orbit coupling, the $x$ component of the kinetic momentum does not commute with the single-particle Hamiltonian (2), unless the Raman coupling $\Omega_{R}$ vanishes. This implies the lack of Galilean invariance in spin-orbit-coupled BECs, which has significant implications for the superfluid behavior, as discussed in the introduction. In the context of this work, the key implication is that constructing states with a constant superflow is not the same as constructing states that propagate at a constant velocity. The former involves fixing the kinetic momentum $\langle\bm{P}\rangle$, as in the calculations of the superfluid fraction in [56, 57] and in the study of current-carrying supersolid configurations in [64]. In contrast, here we focus on moving supersolid patterns with fixed canonical momentum $\langle\bm{p}\rangle$. For clarity we will henceforth refer to this quantity simply as “momentum,” omitting the qualifier “canonical.”
@@
The ground state properties of an interacting spin-orbit-coupled BEC can be understood by first examining the structure of the single-particle energy spectrum. Upon diagonalization of the Hamiltonian (2) at fixed momentum $\bm{p}$, two energy branches emerge. The lower branch exhibits either two degenerate minima at momenta ±$\pm k_{1}^{\mathrm{SP}}\hat{\bm{e}}_{x}=\pm k_{R}\sqrt{1-(\hbar\Omega_{R}/4E_{R})^{2}}\hat{\bm{e}}_{x}$ (for $\hbar\Omega_{R}<4E_{R}$) or a single minimum at zero momentum (when $\hbar\Omega_{R}\geq 4E_{R}$[68]. The interplay between this structure and interaction effects gives rise to a rich equilibrium phase diagram. The phase structure can be inferred by minimizing the total energy (3) as a function of the system parameters [69, 70] (for comprehensive overviews, see also the reviews in Refs. [71, 72, 73, 74, 75]). Among the three quantum phases that emerge, the stripe phase is of particular interest. This phase appears at low Raman coupling in the presence of an antiferromagnetic spin-dependent interaction, i.e., when $g_{ss}>0$. In this regime, both single-particle minima–corresponding to the dressed spin states mentioned in the introduction–are equally populated, resulting in a configuration with zero spin polarization $\langle\sigma_{z}\rangle=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\sigma_{z}\Psi$. This balanced occupation leads to the formation of spatial modulations in the density profile. The emergence of such modulations indicates spontaneous breaking of translation symmetry. When combined with the global $\mathrm{U}(1)$ phase symmetry breaking intrinsic to Bose-Einstein condensation, this reveals the supersolid nature of the stripe phase. This interpretation is further corroborated by the properties of the Bogoliubov spectrum [76, 58] (see also Sec. V).
@@
In addition to the stripe phase, which is the main focus of this paper, the equilibrium phase diagram of spin-orbit-coupled BECs includes two non-supersolid phases [69, 70]. The plane-wave phase corresponds to a scenario where atoms occupy only one of the two single-particle energy minima. Depending on which minimum is chosen, the momentum and spin polarization can take opposite values: $\langle\bm{p}\rangle=\pm N\hbar k_{1}^{\mathrm{PW}}\hat{\bm{e}}_{x}$, $\langle\sigma_{z}\rangle=\pm Nk_{1}^{\mathrm{PW}}/k_{R}$, where
@@
and $\Omega_{\mathrm{cr2}}$ is the critical Raman coupling defined below. In the plane-wave phase, the translation symmetry is unbroken, and the density remains uniform and equal to the average density $\bar{n}=N/V$. This feature is shared by the second non-supersolid phase, the single-minimum phase in which both the momentum and spin polarization are zero.
@@ III Traveling wave patterns
We will now demonstrate how to construct solutions to the Gross-Pitaevskii equation (1) that describe stripe patterns moving with constant velocity. This process will be illustrated in both the laboratory (Sec. III.1) and comoving (Sec. III.2) frames. The laboratory frame corresponds to the rest frame of the Raman lasers responsible for generating the spin-orbit coupling, while the comoving frame is defined as the rest frame of the crystal pattern. We will show that, despite the lack of Galilean invariance in spin-orbit-coupled BECs, these two frames provide equivalent descriptions. Finally, in Sec. III.3, we will explore the key properties of the order parameter associated with the moving stripes.
III.1 Laboratory frame
Consider a periodic wave propagating in a spin-orbit-coupled BEC with a constant velocity $\bm{v}$ in the laboratory frame. The order parameter takes the standard form
$\Psi(\bm{r},t)=\mathrm{e}^{-\mathrm{i}\mu t/\hbar}\Psi_{0}(\bm{r}-\bm{v}t)\,,$ (7)

### v2

@@ Abstract
We consider a traveling supersolid stripe pattern in a spin-orbit-coupled Bose gas. This configuration is associated with an unequal population of the two single-particle energy minima, giving rise to a chemical potential difference that sets the fringe velocity. Unlike stationary stripes, the moving pattern is spin-polarized, with decreasing contrast as the population imbalance increases, eventually leading to stripe melting and transition to the uniform plane-wave phase. The Bogoliubov spectrum of the moving stripes exhibits asymmetry under inversion of the excitation quasimomentum. At high population imbalance, we identify energetic and dynamical instabilities in the spin-phonon mode which transforms to the roton mode of the plane-wave phase as the stripe structure vanishes.
@@ I Introduction
In recent years, the interest in supersolidity, where superfluidity coexists with a crystal-like structure, has grown significantly (see reviews in [1, 2, 3, 4, 5, 6]). First discussed in the 1950s [7, 8, 9] and further developed in subsequent decades [10, 11, 12, 13, 14, 16, 15, 17, 18], the supersolidity has only recently been observed in ultracold boson systems inside optical resonators [19], with spin-orbit coupling [20, 21, 22], and with dipolar interactions [23, 24, 25, 26, 27], as well as in exciton-polariton systems [28, 29]. Patterned states with supersolid-like dynamical features have also been found in driven superfluids [30]. These advances sparked studies on several aspects of supersolids, including their ability to sustain a dissipationless flow, i.e., a supercurrent. At low flow velocities, the superfluid fraction, bounded strictly below one due to crystalline order [13, 18], is inferred from the system’s response to a translation [31, 32, 33, 34, 35, 36, 37] or rotation [40, 38, 41, 39, 42, 43] constraint. However, accessing this quantity in experiments has proven challenging [44, 45], and only very recently a measurement has been performed [46]. At higher flow velocities, supercurrent-carrying supersolids can become unstable [47, 48, 33, 49], and in rotating configurations this can lead to the formation of vortices [17, 50, 51, 52], as confirmed in recent experiments [53, 54]. An anomalous Doppler effect has also been predicted [55]. Interestingly, although these results indicate that supersolids are (under certain conditions) able to sustain a supercurrent, frictionless motion of external objects remains impossible [17, 56].
@@
A common method to generate a supercurrent in a Bose-Einstein condensate (BEC) is to impose a phase twist on its time-independent order parameter [57]. In supersolids at zero temperature, this leads to a configuration where the normal component associated with the lattice remains at rest, while the superfluid background flows at a velocity $\bm{v}_{s}$ determined by the twist angle. An alternative approach involves considering a moving lattice pattern with velocity $\bm{v}_{n}=-\bm{v}_{s}$, making the time-dependent order parameter a function of $\bm{r}-\bm{v}_{n}t$. In Galilean-invariant systems, such as dipolar gases, these two descriptions are physically equivalent, being connected by a Galilean transformation. However, spin-orbit-coupled BECs lack Galilean invariance, a feature that profoundly alters their superfluid behavior even in the non-supersolid phases. Early studies have shown that in these phases the superfluid fraction is strictly less than unity even at zero temperature [58, 59, 60]. In addition, the critical velocity for supercurrent stability differs from the threshold for frictionless impurity motion [61, 62], and current-carrying configurations can exhibit both energetic and dynamical instabilities [63], unlike in standard condensates, where only energetic (Landau) instabilities are present [64, 65]. In the supersolid stripe phase, the combined lack of translational and Galilean invariance further reduces the superfluid density compared to the Leggett bound [60] and leads to peculiar stability conditions for current-carrying states [66].
@@
The Raman lasers responsible for generating spin-orbit coupling play a crucial role in the construction of supercurrent states within the stripe phase. In the analysis of Ref. [66], the current is generated by imposing twisted boundary conditions on the condensate order parameter. This procedure effectively constrains the system’s kinetic momentum and describes a scenario in which the superfluid background flows relative to the laboratory frame (i.e., the rest frame of the Raman lasers), while the density modulations remain stationary, as shown in Fig. 1(a1). This is consistent with the idea that the Raman lasers pin the normal component [58]. In the rest frame of the superfluid, both the fringes and the lasers appear to move at the same velocity [see Fig. 1(a2)]. These configurations correspond to the two contrasting, but physically equivalent, pictures discussed earlier: background flow versus lattice motion.
In this work, we unveil a third scenario, the one in which the density modulations move relative to the laboratory frame, with no net mass transport at the coarse-grained (hydrodynamic) level, as illustrated in Fig. 1(b1). This relative motion renders the configuration physically distinct from the two previously discussed cases. Here, the relevant constrained quantity is the canonical momentum, which, unlike the kinetic momentum, commutes with the spin-orbit Hamiltonian [63].
Figure 1: Schematic illustration of background flow and lattice motion scenarios in the stripe phase of a spin-orbit-coupled BEC. (a1) Supercurrent state in the laboratory frame: the superfluid background (horizontal black line) flows, while the Raman lasers generating the spin-orbit coupling (blue and red arrows) and the density modulations (green wavy line) are at rest. (a2) The same supercurrent state in the rest frame of the superfluid, where the Raman lasers and the fringes translate together. (b1) Traveling stripe pattern in the laboratory frame, characterized by the absence of net mass transport at the coarse-grained level. (b2) Traveling stripe pattern in the comoving frame, where the fringes are at rest (see Sec. V). In all scenarios, the normal component is pinned to the Raman lasers.
The traveling stripe patterns that fall within this third scenario can be understood as a generalization of the notion of an imbalanced BEC mixture to the context of spin-orbit coupling, where the relevant degrees of freedom are dressed spin states rather than bare spin components. Specifically, they emerge in the presence of a population imbalance between the two dressed spins entering the condensate order parameter, that is, the two degenerate minima of the single-particle dispersion. This imbalance, quantified by a dimensionless variable $1\leq s_{p}\leq 1$ [see Eqs. (13) and (14) below for a more precise definition], leads to the appearance of a chemical potential difference between the two states, which, in turn, drives the density fringes to move at a constant velocity. Traveling stripes are therefore intermediate structures that smoothly interpolate between the stationary stripe phase, i.e., a balanced mixture of two dressed spin states ($s_{p}=0$), and the uniform plane-wave phase corresponding to a fully polarized configuration in dressed spin ($s_{p}=\pm 1$). For infinitesimal $s_{p}$ the fringe motion corresponds to the crystal Goldstone mode of the stripe phase and can be excited by releasing a weak spin perturbation [67, 68]. The goal of this work is to systematically study traveling stripe configurations beyond the low-$s_{p}$ regime and highlight their distinctive properties. We find that moving stripes have higher energy and reduced contrast compared to stationary stripes. In addition, they exhibit asymmetric density profiles in the two bare spin components and finite bare spin polarization. All these effects intensify with increasing population imbalance $s_{p}$, till the fringes disappear and the system turns into a uniform plane-wave condensate. Further differences with respect to stationary stripes emerge at the level of dynamics. An analysis of the excitation spectrum of moving stripes reveals its asymmetry with respect to the center of the first Brillouin zone. At sufficiently high $s_{p}$ energetic and dynamical instabilities can emerge, which are not present in the stationary ground-state stripe phase. These instabilities involve the spin phonon branch, that becomes the mode with the roton minimum of the plane-wave phase in the $s_{p}\to\pm 1$ limit [69].
@@
This paper is structured as follows. After briefly reviewing the model for a BEC with Raman-induced spin-orbit coupling (Sec. II), we outline the construction of traveling stripe solutions (Sec. III). These solutions are then derived and analyzed in detail (Sec. IV), with numerical results compared with perturbative analytical estimates. Section V focuses on the Bogoliubov spectrum and the instabilities of the moving stripes. The experimental feasibility of observing these configurations is discussed in Sec. VI. Conclusions are presented in Sec. VII. Technical details of the perturbative approach and full expressions of relevant coefficients appearing in the formulas of the main text are given in Appendices A and B, respectively.
@@ II The model
where $\bm{p}=-\mathrm{i}\hbar\nabla_{\bm{r}}$ is the canonical momentum operator. The spin-orbit coupling strength is proportional to the momentum imparted by the Raman lasers, which is $2\hbar k_{R}\hat{\bm{e}}_{x}$, with $\hat{\bm{e}}_{x}$ the unit vector along the $x$ axis. The corresponding energy scale is $E_{R}=\hbar^{2}k_{R}^{2}/2m$. The parameter $\Omega_{R}$ quantifies the Raman coupling strength, while $\delta_{R}$ represents the Raman detuning, which we shall set to zero from now on.
@@
is conserved. On the other hand, the bare spin polarization, given by
$\langle\sigma_{z}\rangle=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\sigma_{z}\Psi\,$, (6)
is conserved only if $\Omega_{R}=0$, as the Raman coupling term in the single-particle Hamiltonian (2) does not commute with $\sigma_{z}$. We point out that in the thermodynamic limit, where both the particle number and the system volume become infinitely large, the conservation laws of $E$, NN, $\langle\bm{p}\rangle$, and $\langle\sigma_{z}\rangle$ (for $\Omega_{R}=0$) are replaced by those of the corresponding average densities, which remain finite in this limit:
$\bar{\varepsilon}=\frac{E}{V}\,,\quad\bar{n}=\frac{N}{V}\,,\quad\bar{\bm{p}}=\frac{\langle\bm{p}\rangle}{V}\,,\quad\bar{s}_{z}=\frac{\langle\sigma_{z}\rangle}{V}\,.$ (7)
@@
We note that the canonical momentum $\bm{p}$ differs from the kinetic momentum $\bm{P}=\bm{p}-\hbar k_{R}\sigma_{z}\hat{\bm{e}}_{x}$, that is, the operator whose expectation value, $\langle\bm{P}\rangle=\int_{V}\mathrm{d}^{3}r\,\Psi^{\dagger}\bm{P}\Psi$, gives the center-of-mass velocity of the system multiplied by the total mass $Nm$. One can also write $\langle\bm{P}\rangle=m\int_{V}\mathrm{d}^{3}r\,\bm{j}_{d}$, where
$\bm{j}_{d}=\frac{\hbar}{2\mathrm{i}m}\left(\Psi^{\dagger}\nabla_{\bm{r}}\Psi-\Psi^{T}\nabla_{\bm{r}}\Psi^{*}\right)-\frac{\hbar k_{R}}{m}(\Psi^{\dagger}\sigma_{z}\Psi)\hat{\bm{e}}_{x}$ (8)
is the particle current density, that obeys the standard continuity equation [69]
$\partial_{t}n+\nabla_{\bm{r}}\cdot\bm{j}_{d}=0\,.$ (9)
@@
Due to the term $\hbar k_{R}\sigma_{z}$ arising from the spin-orbit coupling, the $x$ component of the kinetic momentum does not commute with the single-particle Hamiltonian (2), unless the Raman coupling $\Omega_{R}$ vanishes. This implies the lack of Galilean invariance in spin-orbit-coupled BECs, which has significant implications for the superfluid behavior, as discussed in the Introduction. In the context of this work, the key implication is that constructing states with a constant superflow is not the same as constructing states that propagate at a constant velocity. The former involves fixing the kinetic momentum $\langle\bm{P}\rangle$, as in the calculations of the superfluid fraction in [58, 59] and in the study of current-carrying supersolid configurations in [66]. In contrast, here we focus on moving supersolid patterns with fixed canonical momentum $\langle\bm{p}\rangle$. For clarity we will henceforth refer to this quantity simply as “momentum,” omitting the qualifier “canonical.”
@@
The ground state properties of an interacting spin-orbit-coupled BEC can be understood by first examining the structure of the single-particle energy spectrum. Upon diagonalization of the Hamiltonian (2) at fixed momentum $\bm{p}$, two energy branches emerge. The lower branch exhibits either two degenerate minima at momenta $\pm k_{1}^{\mathrm{SP}}\hat{\bm{e}}_{x}=\pm k_{R}\sqrt{1-(\hbar\Omega_{R}/4E_{R})^{2}}\hat{\bm{e}}_{x}$ (for $\hbar\Omega_{R}<4E_{R}$) or a single minimum at zero momentum (when $\hbar\Omega_{R}\geq 4E_{R}$[70]. The interplay between this structure and interaction effects gives rise to a rich equilibrium phase diagram. The phase structure can be inferred by minimizing the total energy (3) as a function of the system parameters [71, 72] (for comprehensive overviews, see also the reviews in Refs. [73, 74, 75, 76, 77]). Among the three quantum phases that emerge, the stripe phase is of particular interest. This phase appears at low Raman coupling in the presence of an antiferromagnetic spin-dependent interaction, i.e., when $g_{ss}>0$. In this regime, both single-particle minima, corresponding to the dressed spin states mentioned in the Introduction, are equally populated, resulting in a configuration with zero $\bar{s}_{z}$. This balanced occupation leads to the formation of spatial modulations in the density profile. The emergence of such modulations indicates spontaneous breaking of translation symmetry. When combined with the global $\mathrm{U}(1)$ phase symmetry breaking intrinsic to Bose-Einstein condensation, this reveals the supersolid nature of the stripe phase. This interpretation is further corroborated by the properties of the Bogoliubov spectrum [78, 60] (see also Sec. V).
@@
In addition to the stripe phase, which is the main focus of this paper, the equilibrium phase diagram of spin-orbit-coupled BECs includes two non-supersolid phases [71, 72]. The plane-wave phase corresponds to a scenario where atoms occupy only one of the two single-particle energy minima. Depending on which minimum is chosen, the momentum and bare spin polarization densities can take opposite values: $\bar{\bm{p}}=\pm\bar{n}\hbar k_{1}^{\mathrm{PW}}\hat{\bm{e}}_{x}$, $\bar{s}_{z}=\pm\bar{n}k_{1}^{\mathrm{PW}}/k_{R}$, where
@@
and $\Omega_{\mathrm{cr2}}$ is the critical Raman coupling defined below. In the plane-wave phase, the translation symmetry is unbroken, and the density remains uniform and equal to the average density $\bar{n}$. This feature is shared by the second non-supersolid phase, the single-minimum phase in which both the momentum and spin polarization are zero.
@@ III Traveling wave patterns
We will now demonstrate how to construct solutions to the Gross-Pitaevskii equation (1) that describe stripe patterns moving with constant velocity. We first illustrate the emergence of traveling stripes using an approximate two-harmonic Ansatz for the order parameter that has a transparent physical interpretation (Sec. III.1). We then introduce a Bloch-wave Ansatz that solves Eq. (1) exactly, and explore several of its key properties (Sec. III.2). Finally, in Sec. III.3, we describe the numerical and perturbative methods used to calculate the order parameter of traveling stripes.
III.1 Emergence of traveling stripes: two-harmonic Ansatz
The physical origin of traveling stripes can be understood by considering the following two-harmonic Ansatz for the condensate order parameter, that represents an immediate generalization of the one introduced in earlier work [71, 72] (see also Refs. [80, 81] for analogous proposals in Rashba spin-orbit-coupled systems) to describe the stationary stripe phase:
$\Psi(\bm{r},t)=\tilde{\Psi}_{+1}\mathrm{e}^{\mathrm{i}k_{+}x}\mathrm{e}^{-\mathrm{i}\mu_{+}t/\hbar}+\tilde{\Psi}_{-1}\mathrm{e}^{\mathrm{i}k_{-}x}\mathrm{e}^{-\mathrm{i}\mu_{-}t/\hbar}\,.$ (11)
This Ansatz describes a configuration where the atoms condense in a mixture of two dressed spin states, one with momentum $\hbar k_{+}>0$ and the other with momentum $\hbar k_{-}<0$, both along the $x$ direction. We denote by $\tilde{\Psi}_{\pm 1}$ the two-component spinor part of the wave functions of the two dressed spins. Different from the fully stationary Ansatz of Refs. [71, 72], here we introduce two distinct chemical potentials $\mu_{+}$ and $\mu_{-}$ for the two dressed spin states, so that the respective populations densities $\tilde{\Psi}_{+1}^{\dagger}\tilde{\Psi}_{+1}$ and $\tilde{\Psi}_{-1}^{\dagger}\tilde{\Psi}_{-1}$ are individually conserved. Equivalently, one can impose the conservation of the average particle density,
$\bar{n}=\tilde{\Psi}_{+1}^{\dagger}\tilde{\Psi}_{+1}+\tilde{\Psi}_{-1}^{\dagger}\tilde{\Psi}_{-1}\,$, (12)
and of the dressed spin polarization density,
$\bar{s}=\tilde{\Psi}_{+1}^{\dagger}\tilde{\Psi}_{+1}-\tilde{\Psi}_{-1}^{\dagger}\tilde{\Psi}_{-1}\,.$ (13)
The corresponding Lagrange multipliers are the mean chemical potential $\mu_{d}=(\mu_{+}+\mu_{-})/2$ and the chemical potential semi-difference $\mu_{s}=(\mu_{+}-\mu_{-})/2$. For later convenience we also define the dimensionless population imbalance between the two dressed spins as the ratio
$s_{p}=\frac{\bar{s}}{\bar{n}}\,.$ (14)
Notice that $1\leq s_{p}\leq 1$ and that the extreme values $s_{p}=\pm 1$ correspond to the macroscopic occupation of a single dressed spin state, that is, to a plane-wave configuration. On the other hand, the value $s_{p}=0$ defines a balanced coherent mixture of two dressed spins, which is precisely the stripe phase at equilibrium. As we will see, in our spin-orbit-coupled BEC $s_{p}$ plays a very important role, similar to that of the bare spin imbalance $\bar{s}_{z}/\bar{n}$ in the absence of spin-orbit coupling. The two quantities actually coincide in the limit of zero Raman coupling, where the dressed spin states approach their bare spin counterparts (see Sec. IV.1).
The conservation of $\bar{s}$ follows from the conservation of the average momentum density $\bar{\bm{p}}$. A simple calculation gives
$\bar{p}_{x}=\hbar(k_{c}\bar{n}+k_{1}\bar{s})\,$, (15)
while $\bar{p}_{y}=\bar{p}_{z}=0$ because the order parameter (11) is independent of $y$ and $z$. In Eq. (15) we have introduced the combinations $k_{c}=(k_{+}+k_{-})/2$ and $k_{1}=(k_{+}-k_{-})/2$. They correspond to the (signed) length of the mean condensation wave vector, $\bm{k}_{c}=k_{c}\hat{\bm{e}}_{x}$, and to the half-length of the stripe wave vector, $2\bm{k}_{1}=2k_{1}\hat{\bm{e}}_{x}$. The latter fixes the spatial periodicity, equal to $\pi/k_{1}$, of the density modulations.
The values of the components of $\tilde{\Psi}_{\pm 1}$, as well as those of $k_{\pm}$ and $\mu_{\pm}$, can be determined minimizing the system’s energy. If one imposes $\mu_{+}=\mu_{-}$ in Eq. (11) and only fixes the value of $\bar{n}$, like in Refs. [71, 72], one finds that, for $\delta_{R}=0$ and $g_{\uparrow\uparrow}=g_{\downarrow\downarrow}$, the ground state has either $s_{p}=0$ (in the stripe phase) or $s_{p}=\pm 1$ (in the plane-wave phase). Conversely, if the two chemical potentials are kept distinct, then $s_{p}$ can take any value between 1-1 and $1$. Remarkably, from the form of the order parameter (11), one sees that, if $\mu_{+}\neq\mu_{-}$, the density modulations travel along $x$ at constant velocity proportional to the chemical potential difference:
$v=\frac{\mu_{s}}{\hbar k_{1}}\,.$ (16)
We thus conclude that a population imbalance between the two dressed spin components of the condensate order parameter (11) induces a translation motion of the density fringes. The aim of this paper is to study in detail the properties of these moving stripes.
It is natural to expect that in the $\bar{n}\to 0$ limit the two dressed spin states entering the Ansatz (11) approach the two single-particle energy minima. Since these two minima have opposite momenta and are degenerate in energy one has $k_{+}=-k_{-}=k_{1}^{\mathrm{SP}}$ and $\mu_{+}=\mu_{-}$ irrespective of the value of the population imbalance $s_{p}$. As we will see below, at finite density and for $s_{p}\neq 0$ interactions can induce a mismatch in both the magnitude of the condensation wave vectors and the chemical potentials, thus being at the origin of the stripe motion.
III.2 Traveling Bloch-wave Ansatz
Although the two-harmonic Ansatz (11) is very useful for obtaining a simple qualitative description of the stripe phase, in both stationary and moving configurations, it must be emphasized that it does not represent an exact solution of the time-dependent Gross-Pitaevskii equation (1), unless the nonlinear couplings $g_{dd}$ and $g_{ss}$ are zero. As shown in previous work on stationary stripes [78, 60], interactions cause the presence, in the condensate order parameter, of higher-order harmonic terms. The same is expected to happen in moving stripes. In this paper we will thus construct and make use of an improved Ansatz. For this purpose we look for solutions of Eq. (1) that travel with a constant velocity $v$ along $x$ in the laboratory frame, and that are uniform along $y$ and $z$. Such solutions have the standard form
$\Psi(\bm{r},t)=\mathrm{e}^{-\mathrm{i}\mu t/\hbar}\Psi_{0}(x-vt)\,$, (17)